OpenStudy (anonymous):

A sample of cyanic acid ( HCNO) contains 4.44 * 10 ^22 atoms. How many moles of cyanic acid are in the sample?

sam (.sam.):

\[4.44 \times 10 ^{22}~atoms \times \frac{1~molecule}{4~atoms} \times \frac{1~mole}{6.02 \times 10^{23}~molecule}\] The equation is quite self explanatory.

OpenStudy (lgbasallote):

@.Sam. how did you get 1 molecule = 4 atoms? because there are 4 elements?

sam (.sam.):

yep